OpenStudy (anonymous):

what is the answer to this equation 2-1+6-3+21+???

OpenStudy (texaschic101):

this is not an equation...there is no equal sign

OpenStudy (anonymous):

sorry i meant problem

OpenStudy (anonymous):

nvm i got the answer

OpenStudy (texaschic101):

2 - 1 + 6 - 3 + 21 combine like terms 2 + 6 + 21 = 29 -3 -1 = -4 so we have : 29 - 4 = 25